1. Given the information below, calculate LTM sales as of 9/30/08. ($ in millions) Sales FYE 12/31/07 $750.0 YTD 9/30/07 600.0 YTD 9/30/08 850.0
Answer #1

Sales from 01/01/06 to 12/31/07 = $750 millions

Sales from 01/01/06 to 09/30/07 = $600 millions

So ,sales from 09/30/07 to 12/31/07 = total sales in FY 17 - sales from 01/01/06 to 09/30/07

Sales from 09/30/07 to 12/31/07 = 750 - 600 = $150 millions

Sales from 01/01/08 to 09/30/08 = $850 millions

So, LTM(Last Twelve Months) sales as of 09/30/08 = Sales from 01/01/08 to 09/30/08 + Sales from 09/30/07 to 12/31/07

LTM sales as of 09/30/08 = 850 + 150

LTM sales as of 09/30/08 = $1,000 millions